# FlatFusion: Delving into Details of Sparse Transformer-based Camera-LiDAR Fusion for Autonomous Driving

### v1

@@ Abstract
The integration of data from diverse sensor modalities (e.g., camera and LiDAR) constitutes a prevalent methodology within the ambit of autonomous driving scenarios. Recent advancements in efficient point cloud transformers have underscored the efficacy of integrating information in sparse formats. When it comes to fusion, since image patches are dense in pixel space with ambiguous depth, it necessitates additional design considerations for effective fusion. In this paper, we conduct a comprehensive exploration of design choices for Transformer-based sparse camera-LiDAR fusion. This investigation encompasses strategies for image-to-3D and LiDAR-to-2D mapping, attention neighbor grouping, single modal tokenizer, and micro-structure of Transformer. By amalgamating the most effective principles uncovered through our investigation, we introduce FlatFusion, a carefully designed framework for sparse camera-LiDAR fusion. Notably, FlatFusion significantly outperforms state-of-the-art sparse Transformer-based methods, including UniTR, CMT, and SparseFusion, achieving 73.7 NDS on the nuScenes validation set with 10.1 FPS with PyTorch.
1 Introduction
The pursuit of autonomous driving [14, 68] has necessitated the development of sophisticated 3D object detection systems [59] capable of accurately perceiving and interpreting the environment [21, 22, 32]. At the forefront of this endeavor is the integration of multi-modal sensory inputs [42], with a particular emphasis on the fusion of camera and LiDAR data. The synergy between these sensors—where cameras provide a wealth of semantic information and LiDARs offer precise geometric details, is a compelling avenue for enhancing the robustness and accuracy of autonomous driving systems [22, 23, 20]. Despite the potential, the fusion of camera and LiDAR data presents a formidable challenge due to their distinct characteristics. Cameras capture dense, yet depth-ambiguous [34], 2D images, while LiDAR generates 3D point clouds that, although accurate, are sparse and less semantically rich [75]. This discrepancy demands innovative approaches to integrate their complementary strengths.
@@ 1 Introduction
Recent progress in the domain of sparse Transformers [50, 43, 56] has opened new avenues for efficient data integration in a format that is amenable to the sparsity of LiDAR data, which demonstrates better effectiveness and scalability than SparseConv [7]. However, the dense nature of camera imagery requires careful consideration to enable a seamless fusion with the sparse LiDAR representation. Pioneering works [57, 64, 65] propose the general framework of spare fusion, leaving the specific design of each component unexplored. To fulfill this gap, we delve into the design principles and architectural choices critical for the effective fusion of camera and LiDAR data through the lens of sparse Transformers via in-depth analysis, visualization, and ablations.
@@
2 Related Work
2.1 LiDAR-based 3D Perception
@@
Owing to its capacity for delivering accurate locational data, LiDAR has ascended to a pivotal role as a crucial sensor in ensuring the reliability of autonomous driving systems [61, 24, 19, 10]. Recent LiDAR-centric methods can be broadly categorized into two types according to the form of encoding, i.e., convolution based and Transformer based. More precisely, earlier methods [48, 75, 27] processed point cloud data directly, or transformed it into fixed-size voxels or pillars, then employed 3D CNNs for feature extraction. Additionally, some methods [66, 6, 70, 73, 72] have utilized sparse convolution techniques to further enhance computational efficiency. On the other hand, Transformer has recently become popular in the point cloud field. PCT [12] first computes self-attention globally over the entire point cloud, but its computational complexity grows quadratically with the number of points, making it difficult to adapt to large-scale point clouds. SST [11], DSVT [56] and FlatFormer [43] partition the BEV space into non-overlapping and equally sized windows. These windows are then subjected to a shift similar to that of the SwinTransformer [41] to calculate and exchange window self-attention. Subsequently, PTv3 [62] expands serialized neighbor grouping across multiple tasks.
Figure 2: General Framework of Sparse Transformer-based camera-LiDAR Fusion. The input images and point cloud are fisrt tokenized by their modality-specific backbones. Then, tokens from both modalities are fused in different representation space (3D or pixel space) by projecting tokens from one modality to another. The fusion process is conducted by sparse window attention. Finally, the fused feature are fed into each task-specific head.
2.2 Visual-based 3D Perception
@@
Recently, camera-based 3D perception has gained significant interest due to the cost-effectiveness of cameras [60, 44]. LSS [47] spearheaded the process of transforming image features into 3D space by predicting depth maps from images. Following this, BEVDet [16] and BEVDepth [35] advance this concept, further enhancing the performance of camera-only BEV perception. PETR [40] achieves encoding and decoding by integrating 3D spatial positional embeddings into the image features, leveraging the transformer directly. Unlike these 2D to 3D transformations, BEVFormer [36] utilizes deformable attention [76] to project BEV features onto the image, and computes cross-attention with image features. Recently, several approaches [63, 37, 29] have attempted to combine these two view transformations to achieve more robust BEV representation. Furthermore, diverse temporal fusion strategies [36, 15, 58] have been extensively employed in camera-only BEV perception. Nevertheless, issues such as ambiguity and occlusion of depth remain challenges for pure camera solutions.
2.3 Multi-modal 3D Perception
@@
Combining LiDAR’s accurate position information and RGB camera’s rich semantic information undoubtedly can improve importantly for accurate 3D perception and reliable autonomous driving system [30, 25]. Initial multi-modal approaches [18, 54, 55] were mainly point-based, i.e., using semantic information from images to enhance the point cloud with richer features. The BEV-based fusion methods [42, 38] first encodes BEV features from each modality and then concatenates them. In order to prevent the loss of height information, UVTR [33] employs voxel-level modeling and leverages the transformer decoder to sample features in a uniform space. Transfusion [1] is also based on transformer whose decoder has been widely used in other methods. DeepInteraction [67] proposes a bimodal interaction strategy by alternating modalities of cross attention. As for recent sparse fusion methods: CMT [65], SparseFusion [64], and ObjectFusion [3] align object-centered features across modalities. However, they heavily rely on the feature quality of single branch and thus requires large backbones (e.g., VoVNet, Swin-Tiny) which is less efficient. UniTR [57] uniformly adopts windowed attention to fuse information in different representation spaces, but its projection process is rather inaccurate and thus degenerates performance. GAFusion[31] leverages LiDAR guidance to compensate for depth distribution of the camera features. IS-Fusion [69] jointly captures the Instance- and Scene-level contextual information.
3 Method
@@
(a) Pillar LiDAR Feature Projection [57]
(b) Voxel LiDAR Feature Projection
Figure 3: Comparision of Different 3D-to-2D Fusion Strategies. Pillar LiDAR feature lose $z$-axis information and thus is only projected to limited area in pixel space, compared to the voxel one. The poor correspondence results in degenerated fusion performance.
3.1 General Framework & Experiments Setting
As shown in Fig. 2, sparse Transformer-based camera-LiDAR fusion framework consists of backbones for each modality which tokenize the inputs, two representation spaces to fuse information (2D image plane & 3D space) which requires different projection strategies, and sparse windowed attention blocks which requires different neighbor partition strategies and customized Transformer structure.
In the following sections, all experiments are conducted on nuScenes and results on validation set is reported. Latency is tested in 4090 with batch size 1 and we report the latency of the compared component only. Please refer to Sec. 4 for details. We ensure that all experiments are strictly controlled, allowing only one variable. Final selection and best results in each column are marked.
Table 1: Design of Single Modal Tokenizer
[…]
3.2 Backbones
Raw sensor data, i.e., images and point clouds, must be transformed into tokens before being fed into the Transformer. Given the presence of fusion blocks, it may not be necessary to use overly complex backbones to achieve the optimal balance between performance and efficiency. For image backbones, as shown in Table 1 upper part, we compare: (I) 4 Transforme layers (as in UniTR [57]), (II) ResNet18 (a widely used lightweight encoder), and (III) VoVNet [28] (required by instance query-based methods like CMT [65] and SparseFusion [64]). All image backbones are ImageNet pretrained while 4 Transformer layers are further pretrained in nuImage by [57]. For LiDAR backbones, as shown in Table 1 lower part, we compare: (I) VoxelNet [75] (sparse-conv network), (II) 4 sparse point cloud Transformer layers [56, 43], and (III) None (compensated with 4 additional fusion layers). We conclude:
@@ 3.2 Backbones
Figure 4: 2D-to-3D Projection Strategies. Up to down: Image, Nearest Neighbor, Partial Projection. The projected pixel with larger depth would have a lower intensity in visualization. Nearest Neighbor projection has much worse accuracy of depth compared to Partial Projection. The dropped pixels in Partial Projection are mostly background and thus the information lost is limited.
3.3 Fusion & Projection
@@
Table 2: Design of 3D-to-2D Projection
@@
3D-to-2D: While projecting LiDAR features into pixel space is conceptually straightforward, existing methods such as UniTR [57] encounter limitations. They operate on LiDAR features in the pillar space, which lacks the vertical (z-axis) dimension. Consequently, as depicted in Fig. 3(a), the resulting LiDAR features’ neighbors of image features are confined to the ground plane, leading to a significant loss of contextual information. To address this, we propose to utilize voxel-based LiDAR features for 3D-to-2D fusion, as illustrated in Fig. 3(b). Voxel representations ensures that LiDAR features are comprehensively aligned with the image, providing a fuller representation of the scene. Towards a fair comparison as the voxel representation introduces extra computations, we also experimented with adding 4 additional layers for the pillar baselines, with results summarized in Table 2. Our findings indicate that voxel-based LiDAR features for 3D-to-2D fusion yield substantial performance improvements over the pillar-based representation, despite a moderate increase in computational demand. Notably, augmenting pillar representation with extra layers did not yield performance gains, reinforcing the efficacy of our voxel-based approach.
2D-to-3D: Projecting image features to LiDAR space is complicated by the inherent depth ambiguity in pixels [45]. Therefore, a well-designed projection strategy is required. We investigate three approaches: (I) Nearest Neighbors: projecting according to the nearest voxels in pixel space [71, 57]. While avoiding any learnable modules, this method is highly inaccurate because nearby voxels in 2D space may correspond to very different depths, introducing significant errors, as shown in Fig. 4 (Middle).
Table 3: Design of 2D-to-3D Projection
@@
(II) Depth Estimation: training an additional depth network and using the predicted depth for projection. It can be accurate, but requiring burdensome network [28] to achieve accurate depth prediction. (III) Partial Projection: Unlike existing designs, we propose to only project those image patches that have LiDAR points on, ensuring accurate depth. Table 3 shows that Partial Projection is significantly faster than Depth Estimation, while Nearest Neighbors degrades performance. As shown in Fig. 4 (Lower), since image features are usually downsampled, major part of them are projected and thus the information loss is limited. Besides, all image features are used in 3D-to-2D blocks, which could compensate for loss of some pixels and hence we choose accurate and efficient Partial Projection.
Table 4: Design of Order and Necessity
[…]
(a) Dynamic Set (3D space)
(b) Flatten Window (3D space)
(c) Space-Filling Curve (3D space)
(d) Dynamic Set (2D space)
(e) Flatten Window (2D space)
(f) Space-Filling Curve (2D space)
Figure 5: Comparision of Different Partition Strategies. For 3D space, each grid represents a window in Dynamic Set and Flatten Window. Tokens of the same color within a localized area represent that they are partitioned into the same group to conduct attention. Dynamic Set ensures all elements of the same group are within a single window and thus preserve strong locality while in Flatten Window, elements could be cross multiple windows. Sparce-Filling Curve uses mathematical curves for indexing without considering specific windows, which possesses certain locality yet without any hard constraint and thus is efficient. However, its pattern in pixel space is irregular.
Order & Necessity: We explore the sequence in which 2D-to-3D and 3D-to-2D fusion projections are applied: 2D-to-3D first or 3D-to-2D first. Additionally, we conduct experiments with single fusion projections while maintaining a consistent number of Transformer layers to isolate the impact. Our results, as in Table 4, underscore the importance of both fusion projections, which both contribute unique perspectives into the multi-modal association, aligned with findings in [67, 57]. Notably, 3D-to-2D first outperforms 2D-to-3D first approach by a large margin, which can be attributed to the fact that LiDAR voxel features undergo downsampling between stages [75, 56], whereas the z-axis information, crucial for accurate 3D-to-2D fusion, is lost in the downsampling process. 3D-to-2D first provides more fine-grained LiDAR voxel features in z-axis for 3D-to-2D fusion and thus achieves best performance.
Table 5: Design of Partition Algorithms
[…]

### v2

@@ Abstract
The integration of data from various sensor modalities (e.g. camera and LiDAR) constitutes a prevalent methodology within the ambit of autonomous driving scenarios. Recent advancements in efficient point cloud transformers have underscored the efficacy of integrating information in sparse formats. When it comes to fusion, since image patches are dense in pixel space with ambiguous depth, it necessitates additional design considerations for effective fusion. In this paper, we conduct a comprehensive exploration of design choices for transformer-based sparse camera-LiDAR fusion. This investigation encompasses strategies for image-to-3D and LiDAR-to-2D mapping, attention neighbor grouping, single modal tokenizer, and micro-structure of Transformer. By amalgamating the most effective principles uncovered through our investigation, we introduce FlatFusion, a carefully designed framework for sparse camera-LiDAR fusion. Notably, FlatFusion significantly outperforms state-of-the-art sparse Transformer-based methods, including UniTR, CMT, and SparseFusion, achieving 73.7 NDS on the nuScenes validation set with 10.1 FPS with PyTorch.
I INTRODUCTION
The pursuit of autonomous driving [16, 81, 35, 78, 29, 86, 82, 76, 79] has necessitated the development of sophisticated 3D object detection systems [67, 13] capable of accurately perceiving and interpreting the environment [23, 25, 91]. At the forefront of this endeavor is the integration of multi-modal sensory inputs [49], with a particular emphasis on the fusion of camera and LiDAR data. The synergy between these sensors, where cameras provide a wealth of semantic information and LiDARs offer precise geometric details, is a compelling avenue to improve the robustness and accuracy of autonomous driving systems [25, 26, 22, 68, 24, 31, 11, 77]. Despite the potential, the fusion of camera and LiDAR data presents a formidable challenge due to their distinct characteristics. Cameras capture dense, yet depth-ambiguous [41], 2D images, while LiDAR generates 3D point clouds that, although accurate, are sparse and less semantically rich [93, 90]. This discrepancy demands innovative approaches to integrate their complementary strengths.
Recent progress in the domain of sparse Transformers [58, 50, 64] has opened new avenues for efficient data integration in a format that is compatible with the sparsity of LiDAR data, demonstrating better effectiveness and scalability than SparseConv [8]. However, the dense nature of camera imagery requires careful consideration to enable a seamless fusion with the sparse LiDAR representation. Pioneering works [65, 73, 74] propose the general framework of spare fusion, leaving the specific design of each component unexplored. To fulfill this gap, we delve into the design principles and architectural choices critical for the effective fusion of camera and LiDAR data through the lens of sparse Transformers via in-depth analysis, visualization, and ablations.
@@ I INTRODUCTION
Figure 2: General framework of sparse transformer-based camera-LiDAR fusion. The input images and point cloud are fisrt tokenized by their modality-specific backbones. Then, tokens from both modalities are fused in different representation space (3D or pixel space) by projecting tokens from one modality to another. The fusion process is conducted by sparse window attention. Finally, the fused feature are fed into detection head.
@@
II Related Work
II-A LiDAR-based 3D Perception
@@
Owing to its capacity for delivering accurate locational data, LiDAR has ascended to a pivotal role as a crucial sensor in ensuring the reliability of autonomous driving systems [70, 27, 21]. Recent LiDAR-centric methods can be broadly categorized into two types according to the form of encoding, i.e., convolution based and Transformer based. More precisely, earlier methods [55, 93, 33] processed point cloud data directly, or transformed it into fixed-size voxels or pillars, then employed 3D CNNs for feature extraction. Additionally, some methods [75, 7, 84, 89, 88] have utilized sparse convolution techniques to further enhance computational efficiency. On the other hand, Transformer has recently become popular in the point-cloud field. PCT [14] first computes self-attention globally over the entire point cloud, but its computational complexity grows quadratically with the number of points, making it difficult to adapt to large-scale point clouds. SST [12], DSVT [64] and FlatFormer [50] partition the BEV space into non-overlapping and equally sized windows. These windows are then subjected to a shift similar to that of the SwinTransformer [48] to calculate and exchange window self-attention. Subsequently, PTv3 [71] expands the serialized neighbor grouping across multiple tasks.
II-B Visual-based 3D Perception
@@
Recently, camera-based 3D perception has gained significant interest due to the cost-effectiveness of cameras [69, 51, 28]. LSS [54] spearheaded the process of transforming image features into 3D space by predicting depth maps from images. Following this, BEVDet [18] and BEVDepth [41] advance this concept, further enhancing the performance of camera-only BEV perception. PETR [47] achieves encoding and decoding by integrating 3D spatial positional embeddings into the image features, leveraging the transformer directly. Unlike these 2D to 3D transformations, BEVFormer [42] utilizes deformable attention [94] to project BEV features onto the image and computes cross-attention with image features. Recently, several approaches [72, 43, 36] have attempted to combine these two view transformations to achieve more robust BEV representation. Furthermore, diverse temporal fusion strategies [42, 17, 66] have been extensively employed in camera-only BEV perception. Nevertheless, issues such as ambiguity and occlusion of depth remain challenges for pure camera solutions.
II-C Multi-modal 3D Perception
@@
Combining LiDAR’s accurate position information and RGB camera’s rich semantic information undoubtedly can improve notably for accurate 3D perception and reliable autonomous driving system [37, 30, 87]. Initial multi-modal approaches [20, 62, 63] were mainly point-based, i.e., using semantic information from images to enhance the point cloud with richer features. BEV-based fusion methods [49, 45] first encode BEV features of each modality and then concatenate them. In order to prevent the loss of height information, UVTR [39] employs voxel-level modeling and leverages the transformer decoder to sample features in a uniform space. GAFusion[38] leverages LiDAR guidance to compensate for depth distribution of the camera features. IS-Fusion [83] jointly captures the instance- and scene-level contextual information. Besides BEV-based fusion, there exist other fusion approaches. For instance, Transfusion [1] utilizes image-guided object queries to obtain detection boxes. DeepInteraction [80] proposes a bimodal interaction strategy by alternating modalities of cross attention.
As for recent sparse fusion methods: CMT [74], SparseFusion [73], and ObjectFusion [3] align object-centered features across modalities. However, they heavily rely on the feature quality of single branch and thus require large backbones (e.g., VoVNet, Swin-Tiny) which is less efficient. UniTR [65] uniformly adopts windowed attention to fuse information in different representation spaces, but its projection process is rather inaccurate and thus degenerates performance.
III Method
(a) Pillar LiDAR Projection [65]
(b) Voxel LiDAR Projection
Figure 3: Comparison of different 3D-to-2D fusion strategies. Pillar LiDAR feature lose $z$-axis information and thus is only projected to limited area in pixel space, compared to the voxel one. The poor correspondence results in degenerated fusion performance.
@@ III Method
III-A General Framework & Experiments Setting
As shown in Fig. 2, sparse Transformer-based camera-LiDAR fusion framework consists of backbones for each modality which tokenize the inputs, two representation spaces to fuse information (2D image plane & 3D space) which requires different projection strategies, and sparse windowed attention blocks which require different neighbor partition strategies and customized Transformer structure.
In the following sections, all experiments are conducted on nuScenes and results on validation set is reported. Latency is tested in RTX4090 with batch size 1 and we report the latency of the compared component only. Please refer to Sec. IV for details. We ensure that all experiments are strictly controlled, allowing only one variable. Final selection and best results in each column are marked.
III-B Backbones
Raw sensor data, i.e., images and point clouds, must be transformed into tokens before being fed into the Transformer. Given the presence of fusion blocks, it may not be necessary to use overly complex backbones to achieve the optimal balance between performance and efficiency. For image backbones, as shown in Table I upper part, we compare: (I) 4 Transformer layers (as in UniTR [65]), (II) ResNet18 (a widely used lightweight encoder), and (III) VoVNet [34] (required by instance query-based methods like CMT [74] and SparseFusion [73]). All image backbones are ImageNet pretrained while 4 Transformer layers are further pretrained in nuImage by [65]. For LiDAR backbones, as shown in Table I lower part, we compare: (I) VoxelNet [93] (sparse-conv network), (II) 4 sparse point cloud Transformer layers [64, 50], and (III) None (compensated with 4 additional fusion layers). We conclude:
@@ III-B Backbones
TABLE I: Design of single modal tokenizer
[…]
Image Transformer layers [65] 72.3 68.5 7.9ms
ResNet18 + FPN 73.4 70.2 6.1ms
VoVNet + FPN 73.1 70.1 71.2ms
LiDAR VoxelNet 73.2 70.2 16.9ms
Transformer layers 73.4 70.2 10.8ms
Fusion layers 72.5 69.1 17.1ms
III-C Fusion & Projection
@@
3D-to-2D: While projecting LiDAR features into pixel space is conceptually straightforward, existing methods such as UniTR [65] encounter limitations. They operate on LiDAR features in the pillar space, which lacks the vertical (z-axis) dimension. Consequently, as depicted in Fig. 3(a), the resulting LiDAR features’ neighbors of image features are confined to the ground plane, leading to a significant loss of contextual information. To address this, we propose to utilize voxel-based LiDAR features for 3D-to-2D fusion, as illustrated in Fig. 3(b). Voxel representations ensures that LiDAR features are comprehensively aligned with the image, providing a fuller representation of the scene. Towards a fair comparison as the voxel representation introduces extra computations, we also experimented with adding 4 additional layers for the pillar baselines, with results summarized in Table II. Our findings indicate that voxel-based LiDAR features for 3D-to-2D fusion yield substantial performance improvements over the pillar-based representation, despite a moderate increase in computational demand. Notably, augmenting pillar representation with extra layers did not yield performance gains, reinforcing the efficacy of our voxel-based approach.
Figure 4: Depth map for 2D-to-3D projection with different strategies. Up to down: Raw image, Nearest Neighbor, Partial Projection. The projected pixel with larger depth would have a lower intensity in visualization. Nearest Neighbor projection has much worse accuracy of depth compared to Partial Projection. The dropped pixels in Partial Projection are mostly background and thus the information lost is limited.
TABLE II: Design of projection
@@
3D-to-2D Pillar 72.8 69.5 13.2ms
Voxel 73.4 70.2 16.1ms
Pillar (2$\times$ layers) 72.5 69.4 18.7ms
2D-to-3D Nearest Neighbor 72.9 69.8 7.8ms
Depth Estimation 73.4 70.0 16.1ms
Partial Projection 73.4 70.2 10.7ms
2D-to-3D: Projecting image features to LiDAR space is complicated by the inherent depth ambiguity in pixels [52]. Therefore, a well-designed projection strategy is required. We investigate three approaches: (I) Nearest Neighbors: projecting according to the nearest voxels in pixel space [85, 65]. While avoiding any learnable modules, this method is highly inaccurate because nearby voxels in 2D space may correspond to very different depths, introducing significant errors, as shown in Fig. 4 (Middle). (II) Depth Estimation: training an additional depth network and using the predicted depth for projection. It can be accurate, but requiring burdensome network [34] to achieve accurate depth prediction. (III) Partial Projection: Unlike existing designs, we propose to only project those image patches that have LiDAR points on, ensuring accurate depth. Table II shows that Partial Projection is significantly faster than Depth Estimation, while Nearest Neighbors degrades performance. As shown in Fig. 4 (Lower), since image features are usually downsampled, major part of them are projected and thus the information loss is limited. Besides, all image features are used in 3D-to-2D blocks, which could compensate for loss of some pixels and hence we choose accurate and efficient Partial Projection.
Order & Necessity: We explore the sequence in which 2D-to-3D and 3D-to-2D fusion projections are applied: 2D-to-3D first or 3D-to-2D first. Additionally, we conduct experiments with single fusion projections while maintaining a consistent number of Transformer layers to isolate the impact. Our results, as in Table III, underscore the importance of both fusion projections, which both contribute unique perspectives into the multi-modal association, aligned with findings in [80, 65]. Notably, 3D-to-2D first outperforms 2D-to-3D first approach by a large margin, which can be attributed to the fact that LiDAR voxel features undergo downsampling between stages [93, 64], whereas the z-axis information, crucial for accurate 3D-to-2D fusion, is lost in the downsampling process. 3D-to-2D first provides more fine-grained LiDAR voxel features in z-axis for 3D-to-2D fusion and thus achieves best performance.
TABLE III: Design of order and necessity
@@
3D-to-2D First 73.4 70.2 27.5ms
2D-to-3D First 72.9 69.5 29.2ms
3D-to-2D Only 71.8 67.8 25.4ms
2D-to-3D Only 72.5 69.1 24.6ms
III-D Partition
Table IV gives performance comparisons, leading to the following conclusions:
(a) Dynamic Set
(b) Flatten Window
(c) Space-Filling
(d) Dynamic Set
(e) Flatten Window
(f) Space-Filling
Figure 5: Comparison of different partition strategies. Up to down: 3D space, perspective space. For 3D space, each grid represents a window in Dynamic Set and Flatten Window. Tokens of the same color within a localized area represent that they are partitioned into the same group to conduct attention. Dynamic Set ensures all elements of the same group are within a single window and thus preserve strong locality while in Flatten Window, elements could be cross multiple windows. Space-Filling Curve uses mathematical curves for indexing without considering specific windows, which possesses certain locality yet without any hard constraint and thus is efficient. However, its pattern in perspective space is irregular.
Group Size: One key hyper-parameter of sparse windowed attention is the group size, i.e., the number of tokens per set to conduct attention. We investigate the performance of group size under 3 scales: 40, 80, 160, as in Table IV. We find that 80 is suitable, which is quite different from the conclusion in PTv3 [71], where they find larger group size up to 1024 could consistently bring performance gains. The discrepancy could potentially be attributed to their primary focus on indoor point clouds [9], which are typically denser than their outdoor counterparts [2]. We visualize the partitioning results of different group size in Fig. 6. We can find that when using small group size, the receptive field is limited while large group would leads to bad locality, indicating the necessity of tuning.
III-E Transformer Structure
Transformer [61] consists of attention, FFN, residual link, LayerNorm, and position encoding (PE). Following LLaMA family [60], we adopt FlashAttention [10] and SwiGLU [56]. Existing works like FlatFormer [50] and UniTR [65] both adopt PostNorm [61] while modern architecture [60] usually adopts PreNorm [4]. UniTR proposes to use an extra block-wise residual and post-norm. As for PE, we examine window-level local indices [65] or 3D position encoding. In Table V, we find that PreNorm & 3D PE achieves best performance, aligned with modern findings [60].